Document Sample
ICAPOP_Jan10_Minutes Powered By Docstoc
					          Image, Cloud, Aerosol Product Oversight Panel Meeting # 73

                               Thursday, 28 September 2010
                                   9:30 AM, Room 607


I. Action Items (Turk / Lindsey)

II. Special Discussion Items
Status of GOES-12, -13, and -14

D. Lindsey gave a presentation on the switch from GOES-12 to GOES-13 and preliminary
results from the GOES-14 science test. The presentation is located at:

T. Schmit provided the following additional information:
     Work has begun by vendor ITT to possibly correct the sun light intrusion that occurs with
      GOES-13 during KOZ to prevent the dissemination of corrupted images in GVAR.
     The imager SRF have been updated and implemented into the system reducing the bias to
      about zero.

Oversight Panel Co-Chairs Meeting of December 2009

M. Turk indicated that a result of the Oversight Panel Co-Chairs Meeting was the requirement
that the POPs provide annual reports to the SPSRB. No date is set for ICAPOP’s report but the
ICAPOP co-chairs will be calling upon the members to provide input in the development of this

III. Project Status Updates
The following project status updates were provided offline.

Global Tropical Cyclone Formation Probability Product
(Task 1, provided by and Andrea Schumacher, )

      Currently in development through April 2010 with projected operational date of April
      Critical Design Review in March 2010
      Data collection and initial algorithm development completed July 2009.

ADT v8.1.1
(Task 2, provided by and Chris Velden, )

      Current pre-operational phase due to be completed June 2010 w/ op phase beginning Aug
      V8.1.1 delivered by CIMSS to OSDPD
      Matlab MW recode to Fortran started two weeks ago by Pete Keehn (STAR – Dell Perot)
      CCR1026 approved to establish MW flow to OSDPD ADT sateps systems

AMSU-Based Intensity & Structure Estimation from Aqua & MetOp
(Task 3, provided by John Knaff,

      Project has been delayed by:
       1.      Security exchanges at NCEP and TPC
       2.      TPC is overcommitted, can’t provide access to 64-bit JHT workstation, and this
               project has taken back seat to higher operational priorities
       3.      Access to computer at NHC which has access to BUFR libs/data
      CIRA continues to work with NHC (Alison Krautkramer) to resolve these issues

Updated Volcanic Ash and Cloud Top Height (CLAVR-x)
(Task 4, provided by Mike Pavolonis, )

      Issues encountered when porting code to AIX, delaying OSDPD implementation til late
       Jan, when documentation update is complete
      Anticipate pre-operational implementation late spring/early summer…about 3 months
       behind original schedule
      Scientific performance of automated ash detection, height, and mass loading products has
       exceeded the original expectations
      User can configure warning criteria
      In addition to text message, automatically generated, pre-analyzed false color image with
       product images supplied to user

Miscellaneous CLAVR-x Issues
(Provided by Andy Heidinger, )

      SPSRB requests that involve CLAVR-x and GSIP cloud product generation
       modifications were submitted by A. Heidinger on behalf of Gary Hufford and Stan
       Benjamin. Concern was raised as to whether these requests were submitted properly. See
       New Business below for the discussion which followed on this item.
      A. Heidinger is collecting NCEP's cloud product needs during NPOESS. There is no
       plan not to serve them the data we serve them now. There is also a community-concern
       that two of the NPOESS/VIIRS algorithms are flawed. CLAVR-x could fill this void but
       at a minimum ICAPOP should consider how to make sure NCEP gets what they what.
      Metop-A is now doing HRPT again. A. Heidinger suggests we stop processing FRAC
       data and switch to the Metop-A HRPT over areas of interest as this would ease the
       computational burden greatly (if this is still an issue). We should continue to process
       Metop-A GAC for delivery to NCEP.

GOES-O Readiness
(Task 5)
Data Collection (provided by Tim Schmidt,

      First official GOES-14 images were collected from imager (visible and IR) and sounder
      Improved (4 km) resolution of 13.3 um band required changes to GVAR format. Several
       issues with implement the new GVAR format were discovered, communicated, rectified,
       and verified
      Paired detectors on the higher-res 13.3 um band were inadvertently swapped. Now fixed.
      Image navigation issues have been resolved
      Imager and Sounder data collected for a host of schedules, including rapid scan imagery
      Identified GOES sounder calibration issue with respect to averaging calibration slopes
      Initial IASI inter-calibrations issue with both the imager and sounders
      Various products generated (retrievals, winds, clouds, CSBT, SST, etc.)
      Imagery posted to both the STAR and NESDIS home pages
      More info:


      No update was provided by OSDPD

MODIS Volcanic Ash
(Task 6, provided by

      Vicente reports product is operational and no longer in development phase.
      Recommends removing from ICAPOP projects list

MODIS L2 TPW and Aerosols and MODIS L2 Clouds
(Tasks 7 and 8, provided by )

      After coordination with PAL will make these tasks Inactive and remove from the
       ICAPOP projects list

MetOp Global Multi-Layered Cloud Top Properties
(Task 9, provided by )

      Errors related to ATOVS cloud parameters (CTT, CTP, ECA) reduced by factor of 10
      Cloud Product Error (Water Vapor, Aerosols) less than 20%, with over 60% of the global
       coverage less than 5%
      Errors associated with Cloud Clearing less than 10% for all channels except HIRS/4
       channel 8 (window) and channel 10 (lowest peaking water vapor channel)

Volcanic SO2 & Ash (EOS OMI/AIRS and AVHRR)
(Task 10, provided by )

      Nothing new to report.

IASI Product System
(Task 11, provided by )
      Currently working to get access to AVHRR data and cloud products from OSDPD (POC:
       Gilberto Vicente, Cathy Nichols, Andy Heidinger)
      Attempting to identify JCSDA requirements for the AVHRR/IASI BUFR table (template
       based on MODIS is available) in order to write AVHRR/IASI BUFR program
      Code to bring IASI and AVHRR data together already written and ready because of
       access to test files to develop the software.

Advanced Clear-Sky Processor for Oceans (ACSPO)
(Task 12, provided by )

No update was provided for this meeting. Previous input from September 2009 appears below.

      ACSPO granules are being produced for NOAA-16, 17, 18, 19 and MetOp-A
       1. For NOAA satellites, 1 hour granules are produced from 4km (GAC) AVHRR data
       2. For MetOp, 1hour GAC granules plus 10-minute granules are produced from 1km
           global FRAC data.
      The granules contain a host of layers (in HDF) including SST brightness temperatures,
       albedos, and aerosol optical depths in three channels when available (channel 1, 2, and
       3a). These AODs are calculated only over ocean.
      Other layers in the file include cloud mask, land mask, land proximity, water proximity,
       earth location, solar and satellite geometry, and modeled brightness temperatures from
       the joint center's Community Radiative Transfer Model (CRTM).
      The initial task, Next Gen MetOp Global and Ocean Watch SST: AVHRR Clear-Sky
       Processor for Oceans (ACSPO), has been completed. Emphasis has now shifted to
       enhancements of this ACSPO application and the name of the application has been
       changed to Advanced Clear-Sky Processor for Oceans since the software will be used for
       GOES-R ABI and NPOESS VIIRS instruments as well.

(Task 13, provided by )

      Product running operationally for POES (N18, N19, Metop-A) and DMSP (F16 SSMIS)
      Working on implementation for NPP/ATMS…already running on ATMS proxy data
      Stand ready to extend MiRS to DMSP F18 SSMIS when becomes available
      Ongoing improvements underway focus on hydrometeors and cloud covariance matrix as
       well as adjusting the biases to boost tropical cloud estimates (known to be
       underestimated) w/o degrading mid and high latitude clouds (which seem reasonable)
      Vertical structures of the CLW profiles are being analyzed, and it appears that the
       radiometric signals actually contain some information into that. We are investigating
       further, however.
      Attempting to increase the cloud amounts in the tropics where it appears to be
       underestimated in those latitudes

III. New Business (All)
      Following the update on GOES-14, discussion turned to product validations that should
       have been conducted during the GOES Science Test. The ICAPOP focused on what
       becomes of the product validations that the PALs are expected to conduct and whether
       the POPs should have access to this information. According to the process in place, L.
       Cambardella is to receive updates from the PALs concerning all product validations
       conducted. It was noted that OSDPD did not provide the requested input to the ICAPOP
       for this meeting. Along the same lines, concern was also raised over GOES-13 readiness
       for operations, which led to an action item (see below).

      The ICAPOP took up the issue raised by A. Heidinger concerning the submission of user
       requests by non-users. J. Paquette and T. Schott indicated that while this has been done
       and accepted in the past the better course of action and of least regret would be for the
       user to make the submission. This does not prevent someone else from drawing up the
       request on the user’s behalf to save the user the time, but the user should make the actual

IV. Review of New Action Items (All)
73-1. T. Schott to follow up with the SPIWG on the process of declaring GOES-13 products

73-2. J. Paquette to review user request submitted by A. Heidinger and contact him if there is a

73-3. M. Turk and D. Lindsey to follow up on A. Heidinger’s recommendation that we begin
processing Metop-A HRTP data now that it is flowing instead of the computationally-expensive
FRAC data.

V. Meeting Adjourned at 10:20 am – Next Meeting – Thu, May 13, 2010

Shared By: